A rectangular lawn has a length that is 3 yards greater than the width. The area of the lawn is 108 square yards.
a) Write the polynomial equation for the area of the lawn. Use the variable x to represent the width of the lawn.

Answer by Alan3354(69443)

x = width
L = length
-------
x*L = 108
L = x + 3
x*(x+3) = 108
-----------------
Find a pair of factors of 108 that differ by 3
1*108 NG
2*54 NG
etc